# Perfect Divisibility and Coloring of Some Bull-Free Graphs

### v3

@@ 2 Notations and Preliminary Results
Let $G$ be a connected locally perfect $(P_{6},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the Lemma. For $1\leq i\leq 5$, let
@@
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
@@
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume that $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$, and we are done. So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is locally perfect, we have that $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
@@
Since $G$ is connected, we claim that $u$ has a neighbor in $V(C)$. On the contrary, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}\subseteq M(V(C))$ and $x_{3}\in N(V(C))$. By symmetry, suppose $x_{3}\sim v_{1}$. By Claim 4.1, without loss of generality, we may assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $x_{3}\in W_{1}$, then $\{x_{3},v_{3},v_{4},x_{2},v_{5}\}$ induces a bull; if $x_{3}\in X_{1}$, then there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{1}v_{2}v_{3}$; otherwise, there is an induced $P_{6}=x_{1}x_{2}x_{3}v_{3}v_{4}v_{5}$. All are contradictions. This implies that $N(u)\cap V(C)\neq\mbox{{\rm\O }}$.
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v,u,v_{2},v_{3},v_{4}\}$ induces a bull. If $u\in W_{1}$, then $\{u,v_{3},v_{4},v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
@@
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
@@
Obviously, we have that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{x,y,z,v_{1},v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.2. $\square$
The proof of the following lemma is similar to that of Lemma 4.2.
@@
Let $G$ be a connected locally perfect $(P_{8},C_{5},bull)$-free graph with $\omega(G)\geq 3$. Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 7-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. For $1\leq i\leq 7$, let
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Claim 4.2
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$, and we are done. So, suppose $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$. Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$. Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$. If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, and we are done by Lemma 4.2. So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$, a contradiction. Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{3},v_{4},u,v_{2},v_{6}\}$ induces a bull. So, $u\not\sim v_{6}$. To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$. Moreover, to avoid an induced bull on $\{v_{3},v_{4},u,v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$. Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, then $u\in Y_{1}$ and we are done. So, $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{v_{1},v_{2},u,v_{3},v_{7}\}$, we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},v_{4},u,v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is locally perfect, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{v_{5},v_{6},u,v_{4},v_{7}\}$ induces a bull. So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$. If $u\sim v_{5}$, then there is a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$. Hence, $u$ is anticomplete to $\{v_{5},v_{6}\}$ by symmetry. And thus $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$. This proves Claim 4.2. $\square$
Claim 4.3
For every $u\in N(v)$, we have that $u\in Y$. And for $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. If $N_{C}(u)=\mbox{{\rm\O }}$, since $G$ is connected, there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{3}\cup Y_{1}\cup Z_{1}$ by Claim 4.2; but then there exists an induced $P_{8}=x_{!}x_{2}x_{3}v_{3}v_{4}v_{5}v_{6}v_{7}$. So $u\in N(V(C))$. By symmetry and Claim 4.2, we may assume that $u\in X_{3}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exist an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 4.2. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.3. $\square$
We need the following lemmas.
@@
[17] Let $G$ be a $(P_{6},C_{3})$-free graphs. Then $\chi(G)\leq 4$, and $G$ contains Mycielski-Grötzsch graph if $\chi(G)=4$.
@@
[10] Let $G$ be a $(P_{8},C_{5},C_{3})$-free graphs. Then $\chi(G)\leq 3$.
Now, we can prove Theorems 1.3 and 1.4.
Proof of Theorems 1.3 and 1.4: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly divisible. We will actually prove the following stronger conclusion: every ($P_{8},C_{5}$, bull)-free graph and every ($P_{6}$, bull, $F$)-free graph are perfectly weight divisible. Suppose not, and let $G$ be a ($H$, bull)-free MNWD graph, where $H\in\{P_{6},\{P_{8},C_{5}\}\}$. By Lemma 3.4, we have that $G$ is a connected locally perfect graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect. By Lemmas 4.14.4 and 4.5, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Lemma 3.5, $G[M(v)]$ contains no odd antihole of number of vertices at least 7. By Theorem 1.1, we have that $G[M(v)]$ contains an odd hole. Therefore, by Lemmas 4.2 and 4.3, $G$ has a homogeneous set, which contradicts with Lemma 3.1. This completes the proof of Theorems 1.3 and 1.4. $\square$
5 Proof of Theorem 1.5
In this section, we will prove Theorem 1.5. A class of graphs is hereditary if it is closed under taking induced subgraphs. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
We say a bull-free graph $G$ is basic if neither $G$ nor $\overline{G}$ contains an odd hole with both a vertex which is complete to the odd hole and a vertex which is anticomplete to the odd hole. Chudnovsky and Safra [8] proved the following two lemmas.

### v5

@@ 2 Notations and Preliminary Results
Let $G$ be a connected $(bull,P_{8},C_{5},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a $7$-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{6}v_{7}v_{1}$ be a 7-hole in $G[M(v)]$. For the remainder of the proof, all subscripts are taken modulo 7. Let $i\in[7]$, define
$X_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i}\}\};$
$Y_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+2}\}\};$
$Z_{i}$ $=$ $\{u\in N(V(C))~|~N_{C}(u)=\{v_{i},v_{i+1},v_{i+2}\}\};$
Claim 3.1
$N(V(C))=X\cup Y\cup Z$.
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z$. Let $u\in N(V(C))$. Without loss of generality, suppose that $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done. So we may assume that $u$ has a neighbor in $\{v_{2},v_{3},\cdots,v_{7}\}$.
Suppose that $u$ is anticomplete to $\{v_{2},v_{7}\}$. Then $u$ has a neighbor in $\{v_{3},v_{4},v_{5},v_{6}\}$. Assume that $u$ is anticomplete to $\{v_{3},v_{6}\}$. If $u\sim v_{4}$, then there is a 5-hole $uv_{1}v_{2}v_{3}v_{4}u$, a contradiction. So, $u\not\sim v_{4}$; similarly, $u\not\sim v_{5}$. Therefore, $u$ is not anticomplete to $\{v_{3},v_{6}\}$ and we may by symmetry assume that $u\sim v_{3}$. If $u\sim v_{6}$, then $u$ has a neighbor in $\{v_{4},v_{5}\}$ to avoid a 5-hole $uv_{3}v_{4}v_{5}v_{6}u$; and we may by symmetry assume that $u\sim v_{4}$, but now $\{v_{4},u,v_{3},v_{2},v_{6}\}$ induces a bull. So, $u\not\sim v_{6}$. To avoid a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$, we have that $u\not\sim v_{5}$. Moreover, to avoid an induced bull on $\{u,v_{3},v_{4},v_{2},v_{5}\}$, we have that $u\not\sim v_{4}$. Now, $N(u)\cap V(C)=\{v_{1},v_{3}\}$, and then $u\in Y_{1}$. So, we may assume that $u$ is not anticomplete to $\{v_{2},v_{7}\}$.
Without loss of generality, $u\sim v_{2}$. To avoid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{7}\}$, we may assume that $u\sim v_{3}$. If $u\sim v_{4}$, then $u\sim v_{5}$ to forbid an induced bull on $\{v_{3},u,v_{4},v_{1},v_{5}\}$; similarly, $u\sim v_{7}$. Since $G$ is odd net-free, $u$ is not complete to $V(C)$, and so $u\not\sim v_{6}$; but then $\{v_{4},v_{5},u,v_{2},v_{6}\}$ induces a bull, a contradiction. So $u\not\sim v_{4}$, and similarly, $u\not\sim v_{7}$.
If $u$ is complete to $\{v_{5},v_{6}\}$, then $\{u,v_{5},v_{6},v_{4},v_{7}\}$ induces a bull. So $u$ has at most one neighbor in $\{v_{5},v_{6}\}$. If $u\sim v_{5}$, then there exists a 5-hole $uv_{5}v_{6}v_{7}v_{1}u$. So, $u\not\sim v_{5}$. Similarly, $u\not\sim v_{6}$. Then $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$; this implies $u\in Z_{1}$. This proves Claim 3.1. $\square$
Claim 3.2
If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
Proof. Let $u\in N(v)$ such that $u\notin N(V(C))$. Then $u\in M(V(C))$. Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$. Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$. Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1. If $x_{h-1}\in X_{1}$, then there exists an induced $P_{8}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $x_{h-1}\in Z_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{7},x_{h-2}\}$ induces a bull. Therefore $x_{h-1}\in Y_{1}$.
Given that $h=3$ as otherwise there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$. So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$. Consequently, to avoid an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$. Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$. We claim that $V(S)$ is a homogeneous set of $G$.
Suppose that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$. So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$. Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$. We next show that
$t\not\sim x_{2}.$ (1)
Suppose to the contrary that $t\sim x_{2}$. By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$. Then $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{8}$-free. Hence $t\in Y\cup Z$ by Claim 3.1.
Suppose that $t\not\sim v_{1}$. Moreover, $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{3},v_{7}\}$ as otherwise $\{t,t_{1},x_{2},v_{1},t^{\prime}\}$ induces a bull, where $t^{\prime}\in V(C)\setminus\{v_{1},v_{2},v_{3},v_{7}\}$ is a neighbor of $t$, a contradiction. If $t\sim v_{3}$, then $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$. Both are contradictions. So $t\not\sim v_{3}$ and $t$ is anticomplete to $V(C)\setminus\{v_{2},v_{7}\}$. By Claim 3.1, $t\in Y_{7}$ as $t$ has at least two neighbors in $V(C)$. But then there exists an induced $P_{8}=v_{6}v_{5}v_{4}v_{3}v_{2}tt_{1}t_{2}$. So $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
Now $t\in Y_{1}\cup Z_{1}$ by Claim 3.1. But then $\{t,x_{2},v_{3},v_{4},t_{2}\}$ induces a bull. This proves (1).
By (1), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{2},t_{1},x_{2},t,v_{3}\}$. Similarly, $t\sim v_{1}$. By Claim 3.1, $t\in Y_{1}\cup Z_{1}$. But then there exists an induced $P_{8}=v_{7}v_{6}v_{5}v_{4}v_{3}tt_{1}t_{2}$, a contradiction. Therefore $V(S)$ is a homogeneous set of $G$. This proves Claim 3.2. $\square$
By Claim 3.2, we may assume that
for every $u\in N(v)$, $u\in N(V(C))$. (2)
Claim 3.3
For every $u\in N(v)$, $u\in Y$. And for any $x,y\in N(v)$ with $x\sim y$, $N_{C}(x)=N_{C}(y)$.
Proof. Let $u\in N(v)$. Recall that $G$ is connected. If $N_{C}(u)=\mbox{{\rm\O }}$, then there exists an induced $P_{3}=x_{1}x_{2}x_{3}$ such that $\{x_{1},x_{2}\}$ is anticomplete to $V(C)$ and $x_{3}$ has a neighbor in $V(C)$. We may by symmetry assume that $x_{3}\in X_{1}\cup Y_{1}\cup Z_{1}$ by Claim 3.1; but then there exists an induced $P_{8}=x_{1}x_{2}x_{3}v_{1}v_{7}v_{6}v_{5}v_{4}$. So $u\in N(V(C))$. Without loss of generality, we may assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}$. If $u\in X_{1}$, then there is an induced $P_{8}=vuv_{1}v_{2}v_{3}v_{4}v_{5}v_{6}$. If $u\in Z_{1}$, then $\{v_{2},v_{3},u,v,v_{4}\}$ induces a bull. So $u\in Y_{1}\subseteq Y$.
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup(V(C)\setminus\{v_{1},v_{3}\})$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$. To avoid an induced bull on $\{x,y,v_{1},v_{2},z\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{7}\}$. So $z\in N(V(C))$, but $z\notin X\cup Y\cup Z$, which contradicts with Claim 3.1. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 3.1. $\square$
We now proceed to prove Theorem 1.2.
Proof of Theorem 1.2: On the contrary, let $G$ be a (bull, $P_{8},C_{5}$)-free MNWD graph. By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7. By Lemmas 2.4 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Theorem 1.1, we have that $G[M(v)]$ contains a 7-hole as $G$ is ($P_{8},C_{5}$)-free. Therefore, by Lemma 3.1, $G$ has a homogeneous set, which contradicts Lemma 2.1. This completes the proof of Theorem 1.2. $\square$
4 Proof of Theorem 1.3
In this section, we will prove Theorem 1.3. The following useful lemma is important to our proof, and its proof is similar to the proof of Lemma 3.1.
Lemma 4.1
Let $G$ be a connected $(bull,P_{6},odd~net)$-free graph with $\omega(G)\geq 3$, and let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. If $G[M(v)]$ contains a 5-hole, then $G$ has a homogeneous set.
Proof. Let $C=v_{1}v_{2}\cdots v_{5}v_{1}$ be a 5-hole in $G[M(v)]$. From now on, the subscript is modulo 5 in the proof of the lemma. For $i\in[5]$, let
@@
Proof. It suffices to prove that $N(V(C))\subseteq X\cup Y\cup Z\cup W$. Let $u\in N(V(C))$. Without loss of generality, suppose that $u\sim v_{1}$. If $N(u)\cap V(C)=\{v_{1}\}$, then $u\in X_{1}$ and we are done. So, suppose that $u$ has a neighbor in $\{v_{2},v_{3},v_{4},v_{5}\}$.
@@
To forbid an induced bull on $\{u,v_{1},v_{2},v_{3},v_{5}\}$, $u$ has a neighbor in $\{v_{3},v_{5}\}$. By symmetry, we may assume that $u\sim v_{3}$. If $N(u)\cap V(C)=\{v_{1},v_{2},v_{3}\}$, then $u\in Z_{1}$ and we are done. So, we may suppose that $u$ has a neighbor in $\{v_{4},v_{5}\}$. Since $G$ is odd net-free, we have that $u$ is not complete to $V(C)$. Now, $u\in W_{1}$ if $u\sim v_{4}$ and $u\in W_{5}$ if $u\sim v_{5}$. This proves Claim 4.1. $\square$
We next prove that
Claim 4.2
If there exists $u\in N(v)$ such that $u\notin N(V(C))$, then $G$ has a homogeneous set.
Proof. Let $u\in N(v)$ such that $u\notin N(V(C))$. Then $u\in M(V(C))$. Since $G$ is connected, it follows that there exists an induced $P=x_{1}x_{2}\cdots x_{h}$ connecting $uv$ and $C$, where $h\geq 3$, such that $x_{1}\in\{u,v\}$ and $x_{h}\in V(C)$. Without loss of generality, $x_{1}=u$ and $x_{h}=v_{1}$. Since $x_{h-1}\in N(V(C))$, we may by symmetry assume that $x_{h-1}\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$ by Claim 4.1. If $x_{h-1}\in X_{1}$, then there exists an induced $P_{6}=x_{h-2}x_{h-1}v_{1}v_{2}v_{3}v_{4}$. If $x_{h-1}\in Z_{1}\cup W_{1}$, then $\{v_{2},v_{1},x_{h-1},v_{5},x_{h-2}\}$ induces a bull. Therefore $x_{h-1}\in Y_{1}$.
Given that $h=3$ as otherwise there exists an induced $P_{6}=v_{5}v_{4}v_{3}x_{h-1}x_{h-2}x_{h-3}$. So $x_{2}\sim v_{1}$ and $x_{2}\sim v_{3}$ as $x_{h-1}\in Y_{1}$. Consequently, to avoid an induced $P_{6}=v_{5}v_{4}v_{3}x_{2}x_{1}v$, we have that $v\sim x_{2}$. Let $S$ be the component of $G[N(x_{2})\cap M(V(C))]$ which contains the edge $uv$. We claim that $V(S)$ is a homogeneous set of $G$.
Suppose to the contrary that there exists a vertex $t\in V(G)\setminus V(S)$ such that $t$ is neither complete nor anticomplete to $V(S)$. So there exist two vertices $t_{1},t_{2}\in V(S)$ such that $tt_{1}t_{2}$ is an induced $P_{3}$. Clearly $t\notin V(C)$ as $t_{1}\in M(V(C))$. We next show that
$t\not\sim x_{2}.$ (3)
Suppose to the contrary that $t\sim x_{2}$. By the choice of $S$, we have that $t\in N(V(C))$ as $t\notin V(S)$. It is certain that $t$ has at least two neighbors in $V(C)$ because $t$ has a neighbor in $M(V(C))$ and $G$ is $P_{6}$-free.
Suppose that $t\not\sim v_{1}$. Then $t\not\sim v_{4}$ as otherwise $\{t,t_{1},x_{2},v_{1},v_{4}\}$ induces a bull, a contradiction. Consequently, $t\not\sim v_{3}$ as otherwise $\{t,x_{2},v_{3},t_{2},v_{2}\}$ induces a bull whenever $t\not\sim v_{2}$, and $\{t,v_{2},v_{3},v_{1},v_{4}\}$ induces a bull whenever $t\sim v_{2}$. Both are contradictions. So $t$ is anticomplete to $\{v_{1},v_{3},v_{4}\}$. By Claim 4.1, $t\in Y_{5}$. But now there is an induced $P_{6}=v_{4}v_{3}v_{2}tt_{1}t_{2}$. Therefore $t\sim v_{1}$, and by symmetry $t\sim v_{3}$.
Now, $t\sim v_{2}$ as otherwise $\{t,v_{3},x_{2},v_{2},t_{2}\}$ induces a bull. By Claim 4.1, we may by symmetry assume that $t\in Z_{1}\cup W_{1}$. But then $\{v_{2},v_{1},t,v_{5},t_{1}\}$ induces a bull. This proves (3).
By (3), we have that $t\sim v_{3}$ to avoid a bull on $\{t_{1},t_{2},x_{2},t,v_{3}\}$. Similarly, $t\sim v_{1}$. By Claim 4.1, $t\in Y_{1}\cup W_{3}$. But then $v_{5}v_{4}v_{3}tt_{1}t_{2}$ is an induced $P_{6}$ if $t\in Y_{1}$, and $\{v_{5},v_{1},t,v_{2},t_{1}\}$ induces a bull if $t\in W_{3}$. Both are contradictions. Therefore $V(S)$ is a homogeneous set of $G$. This proves Claim 4.2. $\square$
By Claim 4.2, we may assume that for every $u\in N(v)$,
$u\in N(V(C)).$
@@
By Claim 4.1, without loss of generality, we assume that $u\in X_{1}\cup Y_{1}\cup Z_{1}\cup W_{1}$. If $u\in X_{1}$, then there is an induced $P_{6}=vuv_{1}v_{2}v_{3}v_{4}$. If $u\in Z_{1}$, then $\{v_{2},u,v_{3},v_{4},v\}$ induces a bull. If $u\in W_{1}$, then $\{v_{3},v_{4},u,v,v_{5}\}$ induces a bull. So, $u\in Y_{1}$. This proves (4).
@@
Assume for contradiction that $N_{C}(x)\neq N_{C}(y)$. Without loss of generality, suppose that $N_{C}(x)=\{v_{1},v_{3}\}$ by (4). If $N_{C}(y)=\{v_{2},v_{4}\}$, then $\{v,x,y,v_{1},v_{4}\}$ induces a bull. If $N_{C}(y)=\{v_{3},v_{5}\}$, then $\{v_{3},x,y,v_{2},v_{5}\}$ induces a bull. By symmetry, in all other cases, an induced bull also occurs. Hence, $N_{C}(x)=N_{C}(y)$. This proves (5).
@@
It is certain that $\{v,v_{1},v_{3}\}$ is complete to $V(S)$, and $(N(v)\setminus V(S))\cup\{v_{2},v_{4},v_{5}\}$ is anticomplete to $V(S)$. Suppose that $V(S)$ is not a homogeneous set of $G$. There exists a vertex $z$ which has a neighbor and a non-neighbor in $V(S)$. Clearly, $z\notin N(v)\cup\{v\}\cup V(C)$. Since $S$ is connected, it follows that there exists an edge $xy\in G[S]$ such that $z\sim x$ and $z\not\sim y$.
To avoid an induced bull on $\{y,x,v_{1},z,v_{2}\}$, we have that $z\sim v_{2}$. Similarly, we have that $z$ is complete to $\{v_{2},v_{4},v_{5}\}$. So $z\in N(V(C))$. By Claim 4.1, $z$ has a neighbor in $\{v_{1},v_{3}\}$. By symmetry, we may assume that $z\sim v_{1}$. But then $\{v_{1},x,z,v,v_{4}\}$ induces a bull, a contradiction. Therefore, $V(S)$ is a homogeneous set of $G$. This completes the proof of Lemma 4.1. $\square$
Now we begin to prove Theorem 1.3.
Proof of Theorem 1.3: Since the graph $F$ is non-perfectly divisible, it suffices to prove that every (bull, $P_{6}$, $F$)-free graph are perfectly weight divisible. Suppose not, and let $G$ be a (bull, $P_{6},F$)-free MNWD graph. By Lemma 2.3, we have that $G$ is a connected odd net-free graph, and for every vertex $v\in V(G)$, $G[M(v)]$ is imperfect and contains no odd antihole of number of vertices at least 7. By Lemmas 2.5 and 2.6, we have that $\omega(G)\geq 3$ because $G$ is non-perfectly weight divisible.
Let $v\in V(G)$ such that $v$ belongs to a maximum clique of $G$. By Theorem 1.1, we have that $G[M(v)]$ contains a 5-hole. Therefore, by Lemma 4.1, $G$ has a homogeneous set, which contradicts with Lemma 2.1. This completes the proof of Theorem 1.3. $\square$
5 Proof of Theorems 1.5, 1.6 and 1.7
In this section, we will prove Theorems 1.5-1.7. Substituting a vertex $v$ of a graph $G$ by a graph $H$ is an operation that creates a graph obtained from the disjoint union of $H$ and GvG-v by adding an edge between every vertex of $H$ and every neighbor of $v$ in $G$. Let $\cal C$} be a class of graphs. We denote by ${\cal C}_{s}$ the closure of $\cal C$} under taking substitution.
@@
[…]
@@
[…]
@@
((Theorem 6.1 of [2] )) Let $\cal C$} be a class of graphs. If $\cal C$} is hereditary and polynomially $\chi$-bounded with $\chi\leq\omega^{k}$, then ${\cal C}_{s}$ is polynomially $\chi$-bounded with $\chi\leq\omega^{2k+3}$.